Enter the quadratic equation in standard form. Use the caret (^) to enter exponents. For example, enter x2 as x^2.
2x2 - 12 + 2x = (x - 2)2 + 2
I have yet to learn how to do this..

Answer by rfer(16322)

2x^2-12+2x=x^2-2x-2x+4+2
2x^2+2x-12=x^2-4x+6
2x^2+6x-18=x^2
x^2+6x-18=0